OpenStudy (anonymous):

Integrate 1/(xln(x))

OpenStudy (lgbasallote):

let u = lnx

OpenStudy (lgbasallote):

^hint

OpenStudy (anonymous):

I know that part

OpenStudy (lgbasallote):

so what part dont you know?

OpenStudy (anonymous):

i've been out of university for a year and now i can't remember how to do simple integration. i recall that i need to find du/dx du/dx = 1/x du = 1/x (dx) no idea what to do next

OpenStudy (lgbasallote):

good...so \[\frac{1}{x\ln x} = \frac{du}{u}\]

OpenStudy (apoorvk):